v dv/dx = -9x
write v as a function of x
1
Expert's answer
2020-07-12T17:01:30-0400
"v\space dv/dx=-9x"

a first order differential equation is separable


"v\space dv=-9x\space dx""\int v\space dv=\int -9x\space dx""\frac {v^2}{2}=C_1 -\frac{9x^2}{2}""v=\pm \sqrt {C_2 -9x^2}"

Answer: "v(x)=\pm \sqrt {C -9x^2}"